# LibMOON: A Gradient-based MultiObjective OptimizatioN Library in PyTorch

### v1

@@ 3.4 Multiobjective Bayesian Optimization solvers
When the evaluation of objective functions is costly, multiobjective Bayesian optimization (MOBO) is often the preferred methodology for tackling such challenges. While there are several existing libraries, such as Botorch [48] and HEBO [49], that facilite MOBO, they largely overlook algorithms that leverage decomposition techniques like PSL. To bridge this gap, LibMOON also includs three decomposition-based MOBO algorithms, including PSL-MOBO [25], PSL-DirHVEI [25, 50], and DirHV-EGO [50]. In each iteration, these methods build Gaussian process (GP) models for each objectives and generate a batch of query points for true function evaluations.
PSL-MOBO is an extension of PSL method for expensive MOPs. It optimizes the preference-conditional acquisition function $\alpha(\theta|{\bm{\lambda}})$ over an infinite number of preference vectors to generate a set of promising candidates:
$\min_{\phi}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}(\Delta_{m})}\ \left[\alpha(\theta_{\phi}({\bm{\lambda}})|{\bm{\lambda}})\right],\qquad\theta_{\phi}({\bm{\lambda}}):\Delta_{m}\xrightarrow{}\mathbb{R}^{N}.$ (7)
Table 5: Supported preference-conditional acquisition functions. $p({\bm{y}}|\theta,\mathcal{D})=\mathcal{N}\left(\hat{{\bm{\mu}}}(\theta),\hat{\bm{\sigma}}^{2}(\theta)\right)$ represents the posterior distribution over ${\bm{y}}={\bm{L}}(\theta)$ obtained from GP models, $\beta$ is a parameter to balance exploration and exploitation, ${\bm{z}}^{*}=(z_{1}^{*},\ldots,z^{*}_{m})^{T}$ denotes a utopian point of the given MOP, and $\bm{\xi}=(\xi_{1},\ldots,\xi_{m})^{T}$ is a vector induced by the modified Tchebycheff scalarization.
@@
Tch-LCB $\alpha_{\mathrm{{TchLCB}}}(\theta|{\bm{\lambda}})=\underset{1\leq i\leq m}{\max}\{\lambda_{i}(\hat{\mu}_{i}(\theta)-\beta\hat{\sigma}_{i}(\theta)-z_{i}^{*})\}$
DirHV-EI $\alpha_{\mathrm{{DirHVEI}}}(\theta|{\bm{\lambda}})=\mathbb{E}_{{\bm{y}}\sim p({\bm{y}}|\theta,\mathcal{D})}\left[\prod_{i=1}^{m}[\xi_{i}-y_{i}]_{+}\right]$
@@
4 Snippets of LibMOON examples
For the reader’s convenience, we demonstrate some snippets to show how LibMOON solves synthetic problems efficiently by just a few lines of codes. For example codes solving multitask learning problems, please refer to LABEL:sec:app:empirical. LibMOON can be installed by pip install libmoon.
1from libmoon.solver.gradient.methods import EPOSolver
2from libmoon.util import synthetic_init, get_problem, uniform_pref
3
4problem = get_problem(problem_name=’ZDT1’)
5prefs = uniform_pref(n_prob=5, n_obj = problem.n_obj, clip_eps=1e-2)
6solver = EPOSolver(problem, step_size=1e-2, n_iter=1000, tol=1e-2)
7res = solver.solve(x=synthetic_init(problem, prefs), prefs=prefs)
Listing 1: Finding a size-K (K=5) Pareto solutions with four lines of code in LibMOON.
1from libmoon.solver.psl.core_psl import AggPSLSolver
2from libmoon.util import synthetic_init, get_problem, uniform_pref
3
4problem = get_problem(problem_name=’ZDT1’)
5# agg list [’ls’, ’tche’, ’mtche’, ’pbi’, ... ]
6solver = AggPSLSolver(problem, agg=’ls’)
7model = solver.solve()
8prefs = uniform_pref(n_prob=100, n_obj=problem.n_obj, clip_eps=1e-2)
9eval_y = problem.evaluate(model(Tensor(prefs).cuda()))
Listing 2: Pareto set learning in a problem with three lines of solving problem and two lines of evaluating the results in LibMOON.
Code in Listing 1 demonstrates solving a synthetic problem efficiently with four lines of LibMOON code: 1. Define the problem. 2. Specify the preference vectors. 3. Specify the solver. 4. Obtain the results. The result ‘res’ contains five elements: ‘x_opt’ (optimal Pareto solution), ‘y_opt’ (Pareto objective), ‘x_history’ (decision history), ‘y_history’ (objective history), and ‘hv_history’ (hypervolume history). Codes in Listing 2 show that a Pareto set model can be trained by three lines of codes by assigning a user-specific aggregation function and produce predicted Pareto solutions by two lines of code.
@@ 5 Empirical studies
In this section, we study empirical results of LibMOON. Codes are basically run on a person computer with an i7-10700 CPU, and a 3080 GPU with 10G GPU memory. And we also conduct some analysis regarding to GPU ranging from 3080, 4060, and 4090. The empirical results contain five parts,
@@ 5.1 MOO solvers for synthetic problems
We illustrate the performance of various finite Pareto solution solvers using the VLMOP2 problem, formulated as follows:
$\min_{\theta}f_{1}(\theta)$ $=1-\exp\left(-\sum_{i=1}^{n}\left(\theta_{i}-\frac{1}{\sqrt{n}}\right)^{2}\right),$
$\min_{\theta}f_{2}(\theta)$ $=1-\exp\left(-\sum_{i=1}^{n}\left(\theta_{i}+\frac{1}{\sqrt{n}}\right)^{2}\right).$
@@
    Agg-PBI generates ’exact’ Pareto solutions when the coefficient of $d_{2}$ exceeds a specific value, as guaranteed by PBI [53]. However, this parameter is challenging to tune and is influenced by the Pareto front’s curvature. Additionally, PBI can transform a convex multi-objective optimization problem into a non-convex one, making Agg-PBI less recommended.
@@ 5.2 Pareto set learning on synthetic problems
In this section, we present PSL results using the synthetic VLMOP2 problem as an example. Visualizations of the predicted Pareto solutions are shown in Figure 7, and numerical results are reported in Table 7.
@@
From the table and figure, we have the following key findings:
@@ 5.3 MOO solvers for MTL problems
In this section, we investigate the performance of different finite Pareto solution solvers on the Adult dataset, which is a multitask learning task called fairness classification. For this task, the decision variable $\theta$ is the parameter of a fully-connected neural network, with $|\theta|=28033$. The first objective function is the cross entropy loss, and the second objective is the hyperbolic tangent relaxation of the difference of equality of opportunity (DEO) loss [36][Eq. 6]. The visualization on the Adult problem is shown in Figure 5 and the numerical result is shown in Table 8.
@@ 5.4 Pareto set learning on MTL
    The Pareto Front (PF) of this task is nearly convex, indicating that Local Search (LS) can efficiently find the entire Pareto front. Therefore, it is not necessary to employ complex Multi-Objective Optimization (MOO) methods.
@@ 5.5 MOBO for synthetic and real-world problems
Figure 7: HV curve on MOBO problem. Results are averaged on five random seeds. Reference point to calculate HV : [1.2,1.2].
@@ 5.6 GPU acceleration
Table 10: Running time (minutes) for different platforms on MO-MNIST problem.
@@ Conclusion.
LibMOON was made possible with the valuable feedback and comments from a lot of reseachers. We appreciate the help from Hongzong Li, Xuehai Pan, Zhe Zhao, Meitong Liu, Weiduo Liao, Weiyu Chen, Dake Bu, Yilu Liu, Song Lai, Cheng Gong, Prof. Jingda Deng, Prof. Ke Shang, Prof. Genghui Li, Prof. Zhichao Lu, and Prof. Tao Qin. We sincerely thank them for their early feedback and contributions.
@@ A.2 Aggregation functions
    $g^{\mathrm{cosmos}}(\theta)={\bm{\lambda}}^{\top}{\bm{L}}(\theta)-\mu\frac{{\bm{\lambda}}^{\top}{\bm{L}}(\theta)}{\|{\bm{\lambda}}\|\|{\bm{L}}(\theta)\|},$ (8)
    where $\mu$ is a positive weight factor to align the objective vector ${\bm{L}}(\theta)$ with the preference vector ${\bm{\lambda}}$.
@@
    $g^{\mathrm{LS}}(\theta)=\sum_{i=1}^{m}\lambda_{i}L_{i}(\theta).$ (9)
@@
    $g^{\mathrm{Tche}}(\theta)=\max_{1\leq i\leq m}\left\{\lambda_{i}(L_{i}(\theta)-z_{i})\right\},$ (10)
@@
    $g^{\mathrm{mTche}}(\theta)=\max_{1\leq i\leq m}\left\{\frac{L_{i}(\theta)-z_{i}}{\lambda_{i}}\right\},$ (11)
    where $z$ is a reference point, the same as the point used in the Tchebycheff scalarization function.
@@
    $g^{\mathrm{STche}}(\theta)=\frac{1}{h}\log\sum_{1\leq i\leq m}\exp(h\cdot\lambda_{i}(L_{i}(\theta)-z_{i})).$ (12)
    The Smooth Tchebycheff function uses a relaxed Smooth $\max$ operator. The advantage of this approach is that $g^{\mathrm{STche}}(\theta)$ becomes a Smooth function if each objective function $L_{i}(\theta)$ is Smooth, unlike the non-Smooth $g^{\mathrm{Tche}}(\theta)$. Smooth functions generally have faster convergence compared to non-Smooth ones. Similarly, we can define the Smooth modified Tchebycheff function.
@@
    $g^{\mathrm{PBI}}(\theta)=\underbrace{\frac{1}{\left\lVert{\bm{\lambda}}\right\rVert}\cdot\sum_{i=1}^{m}\lambda_{i}L_{i}(\theta)}_{d_{1}}+\mu\underbrace{\left\lVert{\bm{L}}(\theta)-\frac{d_{1}}{\left\lVert{\bm{\lambda}}\right\rVert}\cdot{\bm{\lambda}}\right\rVert}_{d_{2}},$ (13)
@@
    $g^{\mathrm{pnorm}}(\theta)=\left\lVert{\bm{L}}(\theta)-{\bm{z}}\right\rVert_{p}.$ (14)
@@
    $g^{\mathrm{aasf}}(\theta)=g^{\mathrm{mtche}}(\theta)+\rho g^{\mathrm{LS}}(\theta),$ (15)
@@
Table 12: Notation table
@@
$\theta$ Decision variable of an MOP
$\phi$ Decision variable of a Pareto set model
$m$ Number of objectives
$n$ Number of decision variables
$K$ Number of finite Pareto solutions
$\alpha_{i}$ Coefficients of objective functions
${\bm{\lambda}}$ Preference vector
@@ A.3 Metrics
Metrics can be categorized into two groups. The first group evaluates the quality of a set of solutions ${\mathbb{S}}={y^{(1)},\ldots,y^{(N)}}$, with specific metrics such as IGD and FD relying on the known Pareto front for accuracy.
The second group of metrics assesses the quality of individual solutions $y$ when a preference vector $\lambda$ is provided.
Group 1: Metrics for Sets of Solutions
@@
    Hypervolume (HV) ($\uparrow$) [54]: This metric evaluates both the convergence to the Pareto Front (PF) and the diversity of solutions. A low HV value indicates poor convergence, while high HV values imply better performance. The hypervolume is calculated as the volume dominated by at least one solution in the set ${\mathbb{Y}}$ with respect to a reference point ${\bm{r}}$:
    $\mathrm{HV_{\bm{r}}}=\mathrm{Vol}({{\bm{y}}|\exists{\bm{y}}^{\prime}\in{\mathbb{Y}},{\bm{y}}^{\prime}\preceq{\bm{y}}\preceq{\bm{r}}}).$
@@
    Inverted Generational Distance (IGD) [55]: The IGD measures the average distance between points in a reference set ${\mathbb{Z}}$ and the nearest solutions in the set ${\mathbb{S}}$:
    $\operatorname{IGD}({\mathbb{S}})=\frac{1}{|{\mathbb{Z}}|}\left(\sum_{i=1}^{|{\mathbb{Z}}|}\min_{y^{\prime}\in{\mathbb{S}}}\rho(y^{(i)},y^{\prime})^{2}\right)^{1/2}.$
@@
    Fill Distance (FD) [56]: This metric calculates the covering radius of a set of solutions ${\mathbb{S}}$, defined as the maximum minimum distance from any point in the reference set ${\mathbb{Z}}$ to the nearest solution in ${\mathbb{A}}$:
    $\mathrm{FD}({\mathbb{A}})=\max_{y^{\prime}\in{\mathbb{Z}}}\min_{y\in{\mathbb{A}}}\rho(y,y^{\prime}).$ (16)
@@
    $\mathrm{l}_{\min}=\min_{1<i<j<k}\rho(y^{(i)},y^{(j)})$
    where $\rho$ denotes the Euclidean distance.
@@
    $sl{\min}=-\frac{1}{h\cdot k(k-1)}\log\left(\sum_{1<i<j<N}\exp\left(-h\rho\left(y^{(i)},y^{(j)}\right)\right)\right).$ (17)
@@
    $\mathrm{spacing}=\frac{1}{N}\sum_{i=1}^{N}(d_{i}-\bar{d})^{2},\qquad\bar{d}=\frac{1}{N}\sum_{i=1}^{N}d_{i},\qquad d_{i}=\min_{i\neq j}\rho(y^{(i)},y^{(j)}).$ (18)
@@
    $d_{1}=\frac{\langle y-z,\lambda\rangle}{\left\lVert\lambda\right\rVert},\qquad d_{2}=\left\lVert y-(d_{1}\lambda+z)\right\rVert.$ (20)
@@
    Inner Product (IP): This metric measures the alignment of the objective vector $y$ with the preference vector $\lambda$:
    $\mathrm{IP}=\langle y,\lambda\rangle.$ (21)
@@
Those metrics are summarized in Table 13:
@@
A.4 Example codes solving multitask learning problems
The following code example demonstrates solving a multiobjective fairness classification problem in four steps: defining the hyper-parameter dictionary, setting preference vectors, initializing the multitask optimizer, and solving the problem.
1import torch
2from libmoon.problem.mtl.core.pref_set_mtl import MTL_Pref_Solver, MTL_Set_Solver
3from libmoon.util_mtl.util import get_mtl_prefs
4
5kwargs = {
6 ’batch_size’ : 128,
7 ’lr’ : 1e-3,
8 ’epoch’ : 10,
9 ’solver’ : ’pmgda’,
10 ’dataset_name’ : adult,
11 ’obj_normalization’ : False,
12 ’n_prob’ : 10,
13 ’sigma’ : 0.6,
14 ’h_tol’ : 1e-4,
15 ’device’ : torch.device(’cuda’ if torch.cuda.is_available() else ’cpu’),
16 }
17% For set-based method, e.g., PMTL, MOO-SVGD, HV-Grad, use MTL_Set_Solver(**kwargs)
18mtl_solver = MTL_Pref_Solver(**kwargs)
19pref_mat = get_mtl_prefs(args.dataset_name, kwargs[’n_prob’],
20 obj_normalization=kwargs[’obj_normalization’])
21res, res_history, pref_mat = mtl_solver.solve(pref_mat)
Listing 3: An example of searching for a set of Pareto solutions.
A.5 License and usage of LibMOON
The license used for Adult/Compas/Credit follows Creative Commons Attribution 4.0 International (CC BY 4.0), Database Contents License (DbCL) v1.0, and CC0: Public Domain, respectively.
For academic use of LibMOON, please cite our paper or GitHub. Commercial use requires author permission.

### v2

@@ 3.4 Multiobjective Bayesian optimization solvers
When the evaluation of objective functions is costly, multiobjective Bayesian optimization (MOBO) is often the preferred methodology for tackling such challenges. While there are several existing libraries, such as BoTorch [49] and HEBO [50], that facilities MOBO, they largely overlook algorithms that leverage decomposition techniques like PSL. To bridge this gap, LibMOON also includes three decomposition-based MOBO algorithms, including PSL-MOBO [26], PSL-DirHVEI [26, 51], and DirHV-EGO [51]. In each iteration, these methods build Gaussian process (GP) models for each objectives and generate a batch of query points for true function evaluations.
PSL-MOBO is an extension of PSL method for expensive MOPs. It optimizes the preference-conditional acquisition function $\alpha({\bm{\theta}}|{\bm{\lambda}})$ over an infinite number of preference vectors to generate a set of promising candidates:
$\min_{\bm{\phi}}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}({\bm{\Delta}}_{m})}\left[\alpha({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}})|{\bm{\lambda}})\right],\qquad{\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}):{\bm{\Delta}}_{m}\xrightarrow{}\mathbb{R}^{N}.$ (7)
Table 5: Supported preference-conditional acquisition functions. $p({\bm{y}}|{\bm{\theta}},\mathcal{D})=\mathcal{N}\left(\hat{{\bm{\mu}}}({\bm{\theta}}),\hat{\bm{\sigma}}^{2}({\bm{\theta}})\right)$ represents the posterior distribution over ${\bm{y}}={\bm{L}}({\bm{\theta}})$ obtained from GP models, $\beta$ is a parameter to balance exploration and exploitation, ${\bm{z}}^{*}=(z_{1}^{*},\ldots,z^{*}_{m})^{T}$ denotes a utopian point of the given MOP, and $\bm{\xi}=(\xi_{1},\ldots,\xi_{m})^{T}$ is a vector induced by the modified Tchebycheff scalarization.
@@
Tch-LCB $\alpha_{\mathrm{{TchLCB}}}({\bm{\theta}}|{\bm{\lambda}})=\underset{1\leq i\leq m}{\max}\{\lambda_{i}(\hat{\mu}_{i}({\bm{\theta}})-\beta\hat{\sigma}_{i}({\bm{\theta}})-z_{i}^{*})\}$
DirHV-EI $\alpha_{\mathrm{{DirHVEI}}}({\bm{\theta}}|{\bm{\lambda}})=\mathbb{E}_{{\bm{y}}\sim p({\bm{y}}|{\bm{\theta}},\mathcal{D})}\left[\prod_{i=1}^{m}[\xi_{i}-y_{i}]_{+}\right]$
@@ 4 Empirical studies
In this section, we study empirical results of LibMOON. Codes are basically run on a person computer with an i7-10700 CPU, and a 3080 GPU with 10G GPU memory. And we also conduct some analysis regarding to GPU ranging from 3080, 4060, and 4090. The empirical results contain five parts, MOO solvers for synthetic problems, Pareto set learning solvers for synthetic problems, MOO for MTL problems, Pareto set learning solvers for MTL problems, and GPU accelerations. The report of the first subsections .
@@ 4.1 MOO solvers for synthetic problems
We report the performance of various finite Pareto solution solvers using the VLMOP2 problem, formulated as follows:
[…]
@@
    Agg-PBI generates “exact” Pareto solutions when the coefficient of $d_{2}$ exceeds a specific value, as guaranteed by PBI [54]. However, this parameter is challenging to tune and is influenced by the Pareto front’s curvature. Additionally, PBI can transform a convex multi-objective optimization problem into a non-convex one, making Agg-PBI less recommended.
@@ 4.2 Pareto set learning on synthetic problems
In this section, we present PSL results using the synthetic VLMOP2 problem as an example. Visualizations of the predicted Pareto solutions are shown in Figure 4, and numerical results are reported in Table 7. From the table and figure, we have the following key findings:
@@ 4.3 MOO solvers for MTL problems
In this section, we investigate the performance of different finite Pareto solution solvers on the Adult dataset, which is a multitask learning task called fairness classification. For this task, the decision variable $\theta$ is the parameter of a fully-connected neural network, with $|{\bm{\theta}}|=28033$. The first objective function is the cross entropy loss, and the second objective is the hyperbolic tangent relaxation of the difference of equality of opportunity (DEO) loss [37][Eq. 6]. The visualization on the Adult problem is shown in Figure 5 and the numerical result is shown in Table 8.
@@ 4.4 Pareto set learning on MTL
    The Pareto Front (PF) of this task is nearly convex, indicating that Linear scalarazation (LS) can efficiently find the entire Pareto front. Therefore, it is not necessary to employ complex Multi-Objective Optimization (MOO) methods.
@@ 4.5 MOBO for synthetic and real-world problems
Figure 7: HV curves on MOBO problem. Results are averaged on five random seeds. Reference point to calculate HV : [1.2,1.2].
@@ 4.6 GPU acceleration
Table 10: Running time (minutes) for different platforms on the MO-MNIST problem.
@@ Conclusion.
LibMOON was made possible with the valuable feedback and comments from a lot of researchers. We appreciate the help from Hongzong Li, Xuehai Pan, Zhe Zhao, Meitong Liu, Weiduo Liao, Weiyu Chen, Baijiong Lin, Dake Bu, Yilu Liu, Song Lai, Cheng Gong, Prof. Jingda Deng, Prof. Ke Shang, Prof. Genghui Li, Prof. Zhichao Lu, and Prof. Tao Qin. We sincerely thank them for their early feedback and contributions.
@@ A.2 Aggregation functions
    $g^{\mathrm{cosmos}}_{\bm{\lambda}}({\bm{\theta}})={\bm{\lambda}}^{\top}{\bm{L}}({\bm{\theta}})-\mu\frac{{\bm{\lambda}}^{\top}{\bm{L}}({\bm{\theta}})}{\|{\bm{\lambda}}\|\|{\bm{L}}({\bm{\theta}})\|},$ (8)
    where $\mu$ is a positive weight factor to align the objective vector ${\bm{L}}({\bm{\theta}})$ with the preference vector ${\bm{\lambda}}$.
@@
    $g^{\mathrm{LS}}_{\bm{\lambda}}({\bm{\theta}})=\sum_{i=1}^{m}\lambda_{i}L_{i}({\bm{\theta}}).$ (9)
@@
    $g^{\mathrm{Tche}}_{\bm{\lambda}}({\bm{\theta}})=\max_{1\leq i\leq m}\left\{\lambda_{i}(L_{i}({\bm{\theta}})-z_{i})\right\},$ (10)
@@
    $g^{\mathrm{mTche}}_{\bm{\lambda}}({\bm{\theta}})=\max_{1\leq i\leq m}\left\{\frac{L_{i}({\bm{\theta}})-z_{i}}{\lambda_{i}}\right\},$ (11)
    where ${\bm{z}}$ is a reference point, the same as the point used in the Tchebycheff scalarization function.
@@
    $g^{\mathrm{STche}}_{\bm{\lambda}}({\bm{\theta}})=\frac{1}{h}\log\left(\sum_{i=1}^{m}\exp(h\cdot\lambda_{i}(L_{i}({\bm{\theta}})-z_{i}))\right).$ (12)
    The Smooth Tchebycheff function uses a relaxed Smooth $\max$ operator. The advantage of this approach is that $g^{\mathrm{STche}}_{\bm{\lambda}}({\bm{\theta}})$ becomes a Smooth function if each objective function $L_{i}({\bm{\theta}})$ is Smooth, unlike the non-Smooth $g^{\mathrm{Tche}}_{\bm{\lambda}}({\bm{\theta}})$. Smooth functions generally have faster convergence compared to non-Smooth ones. Similarly, we can define the Smooth modified Tchebycheff function.
@@
    $g^{\mathrm{PBI}}_{\bm{\lambda}}({\bm{\theta}})=\underbrace{\frac{1}{\left\lVert{\bm{\lambda}}\right\rVert}\cdot\sum_{i=1}^{m}\lambda_{i}L_{i}({\bm{\theta}})}_{d_{1}}+\mu\underbrace{\left\lVert{\bm{L}}({\bm{\theta}})-\frac{d_{1}}{\left\lVert{\bm{\lambda}}\right\rVert}\cdot{\bm{\lambda}}\right\rVert}_{d_{2}},$ (13)
@@
    $g^{\mathrm{pnorm}}_{\bm{\lambda}}({\bm{\theta}})=\sqrt[\frac{1}{p}]{\left\lVert{\bm{\lambda}}\odot{\bm{L}}({\bm{\theta}})-{\bm{z}}\right\rVert_{p}}.$ (14)
    The symbol $\odot$ denotes the element-wise product between two vectors.
@@
    $g^{\mathrm{AASF}}_{\bm{\lambda}}({\bm{\theta}})=g^{\mathrm{mTche}}_{\bm{\lambda}}({\bm{\theta}})+\rho g^{\mathrm{LS}}_{\bm{\lambda}}({\bm{\theta}}),$ (15)
@@
Table 12: Notations used in this paper
@@
${\bm{\theta}}$ The decision variable of an MOP.
${\bm{\phi}}$ The decision variable of a Pareto set model.
$m$ Number of objectives.
$n$ Number of decision variables.
$K$ Number of finite Pareto solutions.
$\alpha_{i}$ Coefficients of objective functions.
${\bm{\lambda}}$ A preference vector.
@@ A.3 Metrics
Metrics can be categorized into two groups. The first group evaluates the quality of a set of solutions ${\mathbb{Y}}=\{{\bm{y}}^{(1)},\ldots,{\bm{y}}^{(N)}\}$, with specific metrics such as IGD and FD relying on the known Pareto front for accuracy.
The second group of metrics assesses the quality of individual solutions ${\bm{y}}$ when a preference vector ${\bm{\lambda}}$ is provided.
Group 1: Metrics for a set of solutions
@@
    Hypervolume (HV) ($\uparrow$) [55]: This metric evaluates both the convergence to the Pareto Front (PF) and the diversity of solutions. A low HV value indicates poor convergence, while high HV values imply better performance. The hypervolume is calculated as the volume dominated by at least one solution in the set ${\mathbb{S}}$ with respect to a reference point ${\bm{r}}$:
    $\mathrm{HV}_{\bm{r}}({\mathbb{S}})=\mathrm{Vol}({{\bm{y}}\;|\;\exists{\bm{y}}^{\prime}\in{\mathbb{S}},{\bm{y}}^{\prime}\preceq{\bm{y}}\preceq{\bm{r}}}).$
@@
    Inverted Generational Distance (IGD) [56]: IGD measures the average distance between points in a reference set ${\mathbb{Z}}$ and the nearest solutions in the set ${\mathbb{S}}$:
    $\operatorname{IGD}({\mathbb{S}})=\frac{1}{|{\mathbb{Z}}|}\left(\sum_{i=1}^{|{\mathbb{Z}}|}\min_{{\bm{y}}^{\prime}\in{\mathbb{S}}}\rho({\bm{y}}^{(i)},{\bm{y}}^{\prime})^{2}\right)^{1/2}.$
@@
    Fill Distance (FD) [57]: This metric calculates the covering radius of a set of solutions ${\mathbb{S}}$, defined as the maximum minimum distance from any point in the reference set ${\mathbb{Z}}$ to the nearest solution in ${\mathbb{Z}}$:
    $\mathrm{FD}({\mathbb{S}})=\max_{{\bm{y}}^{\prime}\in{\mathbb{Z}}}\min_{{\bm{y}}\in{\mathbb{S}}}\rho({\bm{y}},{\bm{y}}^{\prime}).$ (16)
@@
    $\mathrm{l}_{\min}=\min_{1\leq i<j\leq N}\rho({\bm{y}}^{(i)},{\bm{y}}^{(j)})$
    where $\rho()$ denotes the Euclidean distance.
@@
    $\mathrm{sl}_{\min}=-\frac{1}{h\cdot k(k-1)}\log\left(\sum_{1\leq i<j\leq N}\exp\left(-h\rho\left({\bm{y}}^{(i)},{\bm{y}}^{(j)}\right)\right)\right).$ (17)
@@
    $\mathrm{spacing}=\frac{1}{N}\sum_{i=1}^{N}(d_{i}-\bar{d})^{2},\qquad\bar{d}=\frac{1}{N}\sum_{i=1}^{N}d_{i},\qquad d_{i}=\min_{1\leq i\neq j\leq N}\rho({\bm{y}}^{(i)},{\bm{y}}^{(j)}).$ (18)
@@
    $d_{1}=\frac{\langle{\bm{y}}-{\bm{z}},{\bm{\lambda}}\rangle}{\left\lVert{\bm{\lambda}}\right\rVert},\qquad d_{2}=\left\lVert{\bm{y}}-(d_{1}{\bm{\lambda}}+{\bm{z}})\right\rVert.$ (20)
@@
    Inner Product (IP): This metric measures the alignment of the objective vector ${\bm{y}}$ with the preference vector ${\bm{\lambda}}$:
    $\mathrm{IP}=\langle{\bm{y}},{\bm{\lambda}}\rangle.$ (21)
@@
Those metrics are summarized in Table 13 and also can be found in https://libmoondocs.readthedocs.io/en/latest/apis/Libmoon_metric.html.
@@
A.4 License and usage of LibMOON
The license used for Adult/Compas/Credit follows Creative Commons Attribution 4.0 International (CC BY 4.0), Database Contents License (DbCL) v1.0, and CC0: Public Domain, respectively. For academic use of LibMOON, please cite our paper or GitHub. Commercial use requires author permission.